OpenStudy (anonymous):

Help Please. :c 1.) Whats the inverse of f if f (x) = ^3sqrt x-2 ???? A. f^-1(x)= x^3+8 B. f^-1(x)= 2- ^3sqrt x C. f^-1(x)= x^3+2 D. f^-1(x)= 1 / ^3sqrt x-2 2. What is the simplest form of the expression?? sqrt 2 - sqrt 6 / sqrt 2 + sqrt 6 A. 8-4 sqrt 3 / 8 B. 8+4 sqrt 3 / 8 C. 2 - sqrt 3 D. -2 + sqrt 3

OpenStudy (loser66):

\(f(x) = \sqrt[3] x -2\) or \(f(x) = \sqrt[3]{x-2}\)??

OpenStudy (anonymous):

2nd one.

OpenStudy (anonymous):

What do you think #2 is?

OpenStudy (anonymous):

@DolphinDawg I have no idea , but if I had to guess id say C? idk

OpenStudy (anonymous):

Hmm give me a secnd

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Heres my work : 1 Rationalize the denominator: 6√2√(2√2√)=6√2√2 2√−6√2√2+6√ 2 Simplify 6√2√ to 12−−√ 2√−12−−√2+6√ 3 Simplify 12−−√ to 23√ 2√−23√2+6√ 4 Simplify 23√2 to 3√ 2√−3√+6√ Done Decimal Form: 2.13165249759

OpenStudy (anonymous):

well i need to know for sure D:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

thankyou.. can u help with other one?

OpenStudy (anonymous):

i dont know, i havent turned it in

OpenStudy (anonymous):

yeah & c for both???????????????????

OpenStudy (anonymous):

whats #1

OpenStudy (anonymous):

it was D @DolphinDawg -_-